OpenStudy (anonymous):

A plane is flying east at 115 m/s.the wind accelerates it at 2.88 m/s^2 directly northeast. after 25.0 s what is Tue velocity (magnitude&direction of the plane ?

OpenStudy (anonymous):

if your looking for the final velocity.... (initial) v1 = 115 m/s a = 2.88 m/s^2 t = 25s v = v1 + at V = 115 m/s + 2.88 m/s^2(25s) V = 115 m/s + 72 m/s V = 187 m/s

OpenStudy (anonymous):

yeah thats what i got too but im looking for the magnitude and direction and thas where i have trouble